# Channel Capacity of Near-Field Multiuser Communications

### v1

@@ III-A Sum-Rate Capacity
III-B Capacity Region
@@
${\mathsf{R}}_{1}\leq\log_{2}(1+{P_{1}}{\sigma^{-2}}G_{1}),{\mathsf{R}}_{2}\leq\log_{2}(1+{P_{2}}{\sigma^{-2}}G_{2}),$ (21a)
@@
where ${\mathsf{R}}_{k}$ denotes the achievable rate of UT $k=1,2$. The MAC capacity region forms a pentagon, with its corner points attained through SIC decoding, and the line segment connecting these points achieved through time sharing, as seen in [7, Figure 7]. The corner points are computed as follows.
Lemma 2.
When the SIC decoding order $1\rightarrow 2$ is adopted, the rates of the two UTs are, respectively, given by
${\mathsf{R}}_{1}^{1\rightarrow 2}=\log_{2}\!\left(1+\frac{P_{1}\sigma^{-2}G_{1}+P_{1}P_{2}\sigma^{-4}G_{1}G_{2}\bar{\rho}}{1+P_{2}\sigma^{-2}G_{2}}\right),$ (22a)
${\mathsf{R}}_{2}^{1\rightarrow 2}=\log_{2}\left(1+P_{2}\sigma^{-2}G_{2}\right).$ (22b)
For the decoding order $2\rightarrow 1$, we have
${\mathsf{R}}_{1}^{2\rightarrow 1}=\log_{2}\left(1+P_{1}\sigma^{-2}G_{1}\right),$ (23a)
${\mathsf{R}}_{2}^{2\rightarrow 1}=\log_{2}\left(1+\frac{P_{2}\sigma^{-2}G_{2}+P_{1}P_{2}\sigma^{-4}G_{1}G_{2}\bar{\rho}}{1+P_{1}\sigma^{-2}G_{1}}\right).$ (23b)
Proof:
Please refer to Appendix -D for more details. ∎
Regarding time sharing, it means that applying the SIC order $1\rightarrow 2$ with probability $\tau$, while applying $2\rightarrow 1$ with probability $1-\tau$, where $\tau\in(0,1)$. Given $\tau$, by denoting the achievable rate pair as $({\mathsf{R}}_{1}^{\tau},{\mathsf{R}}_{2}^{\tau})$, we have ${\mathsf{R}}_{1}^{\tau}=\tau{\mathsf{R}}_{1}^{1\rightarrow 2}+(1-\tau){\mathsf{R}}_{1}^{2\rightarrow 1}$ and ${\mathsf{R}}_{2}^{\tau}=\tau{\mathsf{R}}_{2}^{1\rightarrow 2}+(1-\tau){\mathsf{R}}_{2}^{2\rightarrow 1}$. As a result, the MAC capacity region is given by
@@
When $M_{x},M_{z}\rightarrow\infty$, the corner points of the MAC capacity region satisfy
$\begin{split}&\lim_{M_{x},M_{z}\rightarrow\infty}(\mathsf{R}_{1}^{1\rightarrow 2},\mathsf{R}_{2}^{1\rightarrow 2})\approx\lim_{M_{x},M_{z}\rightarrow\infty}(\mathsf{R}_{1}^{2\rightarrow 1},\mathsf{R}_{2}^{2\rightarrow 1})\\ &~{}~{}~{}\approx\left(\log_{2}(1+{P_{1}\xi}/{(2\sigma^{2})}),\log_{2}(1+{P_{2}\xi}/{(2\sigma^{2})})\right).\end{split}$ (25)
@@
III-C Comparison with the FF Capacity
We now focus on the FF case for comparison.
Proposition 1.
Under the FF model, the channel gains are given by $G_{k}^{\mathsf{F}}=\frac{MA\Psi_{k}}{4\pi r_{k}^{2}}$ for $k=1,2$, and the CCF satisfies
$\rho_{{}_{\mathsf{F}}}=\left\{\!\begin{array}[]{ll}1&\left(\theta_{1},\phi_{1}\right)\!=\!\left(\theta_{2},\phi_{2}\right)\\ \frac{1-\cos\left(M_{x}\Delta_{\Phi}\right)}{M^{2}\left(1-\cos\Delta_{\Phi}\right)}&\Phi_{1}\!\neq\!\Phi_{2},\Omega_{1}\!=\!\Omega_{2}\\ \frac{1-\cos\left(M_{z}\Delta_{\Omega}\right)}{M^{2}\left(1-\cos\Delta_{\Omega}\right)}&\Phi_{1}\!=\!\Phi_{2},\Omega_{1}\!\neq\!\Omega_{2}\\ \frac{4\left(1-\cos\left(M_{x}\Delta_{\Phi}\right)\right)\left(1-\cos\left(M_{z}\Delta_{\Omega}\right)\right)}{M^{2}\left(1-\cos\Delta_{\Phi}\right)\left(1-\cos\Delta_{\Omega}\right)}&\text{else}\\ \end{array}\right.,$ (30)
@@ III-C Comparison with the FF Capacity
Please refer to Appendix -E. ∎
Using Proposition 1, the closed-form expressions for the FF MAC capacity and the capacity region follow immediately, which are omitted due to space limitations.
@@
Under the FF model, when $M\rightarrow\infty$, the asymptotic MAC capacity is given in (31), shown at the bottom of this page, which yields $\mathsf{C}_{\mathsf{MAC}}^{\mathsf{F}}\simeq\mathcal{O}\left(\log M\right)$.
@@
The results can be obtained using the fact that $\lim_{M\rightarrow\infty}\rho_{{}_{\mathsf{F}}}=\begin{cases}1&\left(\theta_{1},\phi_{1}\right)=\left(\theta_{2},\phi_{2}\right)\\ 0&\left(\theta_{1},\phi_{1}\right)\neq\left(\theta_{2},\phi_{2}\right)\\ \end{cases}$, along with the approximation $\log_{2}(1+x)\approx\log_{2}x$ for large $x$. ∎
$\mathsf{C}_{\mathsf{MAC}}^{\mathsf{F}}\simeq\begin{cases}\log_{2}\left(\frac{MA}{4\pi\sigma^{2}}\left(\frac{P_{1}\Psi_{1}}{r_{1}^{2}}+\frac{P_{2}\Psi_{2}}{r_{2}^{2}}\right)\right)\triangleq\mathsf{F}_{\mathsf{MAC}}^{\mathsf{s}}&\left(\theta_{1},\phi_{1}\right)=\left(\theta_{2},\phi_{2}\right)\\ \log_{2}\left(\frac{MA}{4\pi\sigma^{2}}\left(\frac{P_{1}\Psi_{1}}{r_{1}^{2}}+\frac{P_{2}\Psi_{2}}{r_{2}^{2}}\right)+\frac{M^{2}A^{2}P_{1}P_{2}\Psi_{1}\Psi_{2}}{16\pi^{2}r_{1}^{2}r_{2}^{2}\sigma^{4}}\right)\triangleq\mathsf{F}_{\mathsf{MAC}}^{\mathsf{d}}&\left(\theta_{1},\phi_{1}\right)\neq\left(\theta_{2},\phi_{2}\right)\\ \end{cases}$ (31)
$\mathsf{F}_{\mathsf{MAC}}^{\mathsf{d}}-\mathsf{F}_{\mathsf{MAC}}^{\mathsf{s}}=\log_{2}\left(1+\frac{MA}{4\pi\left(P_{1}^{-1}\Psi_{1}^{-1}r_{1}^{2}+P_{2}^{-1}\Psi_{2}^{-1}r_{2}^{2}\right)\sigma^{2}}\right)>0$ (32)
@@
This contradiction stems from the assumption of uniform channel power across each antenna element, which becomes increasingly inaccurate as the number of the elements grows. Furthermore, comparing the FF MAC capacity for co-directional UTs and UTs with differing directions, as indicated in (32) below, yields the following observation.
@@
The FF MAC capacity significantly diminishes when UTs are oriented in the same direction compared to cases where they face different directions. This contrast with the NF model arises from the high channel correlation ($\rho_{{}_{\mathsf{F}}}=1$) among UTs in the same direction, leading to notable IUI.
By letting $M\rightarrow\infty$, we characterize the FF MAC capacity region as follows.
Corollary 5.
Under the FF model, when $M\rightarrow\infty$, if UTs are in the same direction, the corner points approximately satisfy $(\mathsf{R}_{1}^{1\rightarrow 2},\mathsf{R}_{2}^{1\rightarrow 2})\simeq(\log_{2}(1+\frac{P_{1}r_{2}^{2}}{P_{2}r_{1}^{2}}),\mathcal{O}(\log M))$ and $(\mathsf{R}_{1}^{2\rightarrow 1},\mathsf{R}_{2}^{2\rightarrow 1})\simeq(\mathcal{O}(\log M),\log_{2}(1+\frac{P_{2}r_{1}^{2}}{P_{1}r_{2}^{2}}))$. If UTs are positioned in different directions, $(\mathsf{R}_{1}^{1\rightarrow 2},\mathsf{R}_{2}^{1\rightarrow 2})\simeq(\mathsf{R}_{1}^{2\rightarrow 1},\mathsf{R}_{2}^{2\rightarrow 1})\simeq\left(\mathcal{O}\left(\log M\right),\mathcal{O}\left(\log M\right)\right)$.
Proof:
Similar to the proof of Corollary 4. ∎
Remark 8.
The results in Corollary 5 indicate that the MAC capacity region under the FF model can expand unboundedly with the number of array elements, rendering it impractical.
Remark 9.
When UTs face different directions, the FF MAC capacity region roughly forms a rectangle and is more extensive than that when UTs are in the same direction.
By comparing the NF capacity with its FF counterpart, we draw the following conclusions:
  • Different from the FF MAC capacity that degenerates for co-directional UTs due to the significant IUI, the NF capacity can be preserved when UTs are located at different locations (different directions or distances), which underscores superior flexibility and robust interference management capabilities inherent to NFC. This additional resolution in the range domain enable space-division multiple access [20] from FF angle-division multiple access to NF range-division multiple access, demonstrating the NFC’s potential as a promising approach for efficient and interference-free communication systems.

### v2

@@ III-A1 Sum-Rate Capacity
We note from (23) that, different from the asymptotic capacity for the UPA given in (21), which is independent of the UTs’ locations, the MAC capacity for the ULA is always affected by the locations of the UTs even as the antenna number approaches infinity. Specifically, it depends on the projected aperture $A\sin\phi_{k}$ and the projected distance to the ULA $r_{k}\sin\theta_{k}$ for each UT.
$\mathsf{C}_{\mathsf{MAC}}\simeq\begin{cases}\log_{2}\left(\frac{MA}{4\pi}\left(\frac{\gamma_{1}\Psi_{1}}{r_{1}^{2}}+\frac{\gamma_{2}\Psi_{2}}{r_{2}^{2}}\right)\right)\triangleq\mathsf{F}_{\mathsf{MAC}}^{\mathsf{s}}&\left(\theta_{1},\phi_{1}\right)=\left(\theta_{2},\phi_{2}\right)\\ \log_{2}\left(\frac{MA}{4\pi}\left(\frac{\gamma_{1}\Psi_{1}}{r_{1}^{2}}+\frac{\gamma_{2}\Psi_{2}}{r_{2}^{2}}\right)+\frac{M^{2}A^{2}\gamma_{1}\gamma_{2}\Psi_{1}\Psi_{2}}{16\pi^{2}r_{1}^{2}r_{2}^{2}}\right)\triangleq\mathsf{F}_{\mathsf{MAC}}^{\mathsf{d}}&\left(\theta_{1},\phi_{1}\right)\neq\left(\theta_{2},\phi_{2}\right)\\ \end{cases}$ (28)
$\mathsf{F}_{\mathsf{MAC}}^{\mathsf{d}}-\mathsf{F}_{\mathsf{MAC}}^{\mathsf{s}}=\log_{2}\left(1+\frac{MA}{4\pi\left(\gamma_{1}^{-1}\Psi_{1}^{-1}r_{1}^{2}+\gamma_{2}^{-1}\Psi_{2}^{-1}r_{2}^{2}\right)}\right)>0$ (29)
III-A2 Capacity Region
@@
${\mathsf{R}}_{1}\leq\log_{2}(1+\gamma_{1}\mathsf{g}_{1}^{\mathsf{n}}),\ {\mathsf{R}}_{2}\leq\log_{2}(1+{\gamma_{2}}\mathsf{g}_{2}^{\mathsf{n}}),$ (24a)
@@
The MAC capacity region forms a pentagon, with two corner points attained through SIC decoding, i.e., $({\mathsf{R}}_{1}^{1\rightarrow 2},{\mathsf{R}}_{2}^{1\rightarrow 2})$ and $({\mathsf{R}}_{1}^{2\rightarrow 1},{\mathsf{R}}_{2}^{2\rightarrow 1})$, and the line segment connecting these points achieved through time sharing, as seen in [11, Figure 7].
Time sharing means that applying the SIC order $1\rightarrow 2$ with probability $\tau$, while applying $2\rightarrow 1$ with probability $1-\tau$, where $\tau\in(0,1)$. Given $\tau$, by denoting the achievable rate pair as $({\mathsf{R}}_{1}^{\tau},{\mathsf{R}}_{2}^{\tau})$, we have ${\mathsf{R}}_{1}^{\tau}=\tau{\mathsf{R}}_{1}^{1\rightarrow 2}+(1-\tau){\mathsf{R}}_{1}^{2\rightarrow 1}$ and ${\mathsf{R}}_{2}^{\tau}=\tau{\mathsf{R}}_{2}^{1\rightarrow 2}+(1-\tau){\mathsf{R}}_{2}^{2\rightarrow 1}$. As a result, the MAC capacity region is given by
@@
When $M_{x},M_{z}\rightarrow\infty$, the corner points of the NF MAC capacity region satisfy
$\lim\nolimits_{M_{x},M_{z}\rightarrow\infty}(\mathsf{R}_{1}^{1\rightarrow 2},\mathsf{R}_{2}^{1\rightarrow 2})\approx\lim\nolimits_{M_{x},M_{z}\rightarrow\infty}(\mathsf{R}_{1}^{2\rightarrow 1},\mathsf{R}_{2}^{2\rightarrow 1})$
$\approx\left(\log_{2}(1+{\xi\gamma_{1}}/2),\log_{2}(1+{\xi\gamma_{2}}/{2})\right).$ (26)
@@
III-B Comparison with the FF Capacity
For comparison, we next discuss the FF case.
Lemma 2.
Under the FF model, the channel gains are given by $\mathsf{g}_{k}^{\mathsf{f}}=\frac{MA\Psi_{k}}{4\pi r_{k}^{2}}$ for $k=1,2$, and the CCF satisfies
$\rho_{\mathsf{f}}=\left\{\!\begin{array}[]{ll}1&\left(\theta_{1},\phi_{1}\right)\!=\!\left(\theta_{2},\phi_{2}\right)\\ \frac{1-\cos\left(M_{x}\Delta_{\Phi}\right)}{M^{2}\left(1-\cos\Delta_{\Phi}\right)}&\Phi_{1}\!\neq\!\Phi_{2},\Omega_{1}\!=\!\Omega_{2}\\ \frac{1-\cos\left(M_{z}\Delta_{\Omega}\right)}{M^{2}\left(1-\cos\Delta_{\Omega}\right)}&\Phi_{1}\!=\!\Phi_{2},\Omega_{1}\!\neq\!\Omega_{2}\\ \frac{4\left(1-\cos\left(M_{x}\Delta_{\Phi}\right)\right)\left(1-\cos\left(M_{z}\Delta_{\Omega}\right)\right)}{M^{2}\left(1-\cos\Delta_{\Phi}\right)\left(1-\cos\Delta_{\Omega}\right)}&\text{else}\\ \end{array}\right.,$ (27)
@@ III-B Comparison with the FF Capacity
The results can be derived from (6) by using the sum of the geometric series and trigonometric identities. ∎
Using Lemma 2, the closed-form expressions for the FF MAC capacity and the capacity region follow immediately, which are omitted due to space limitations.
@@
Under the FF model, when $M\rightarrow\infty$, the asymptotic MAC capacity is given in (28), shown at the bottom of this page, which yields $\mathsf{C}_{\mathsf{MAC}}\simeq\mathcal{O}\left(\log M\right)$.
@@
The results can be obtained using the fact that $\lim_{M\rightarrow\infty}\rho_{\mathsf{f}}=\begin{cases}1&\left(\theta_{1},\phi_{1}\right)=\left(\theta_{2},\phi_{2}\right)\\ 0&\left(\theta_{1},\phi_{1}\right)\neq\left(\theta_{2},\phi_{2}\right)\\ \end{cases}$, along with the approximation $\log_{2}(1+x)\approx\log_{2}x$ for large $x$. ∎
@@
This contradiction stems from the assumption of uniform channel power across each antenna element, which becomes increasingly inaccurate as the number of the elements grows. Furthermore, comparing the FF MAC capacity for co-directional UTs and UTs with differing directions, as indicated in (29), we find that the FF MAC capacity significantly diminishes when UTs are oriented in the same direction. This contrast with the NF model arises from the high channel correlation ($\rho_{\mathsf{f}}=1$) among UTs in the same direction, leading to notable IUI.
We then characterize the asymptotic MAC capacity region under the FF model. When $M\rightarrow\infty$, if UTs are in the same direction, the corner points of the MAC capacity region satisfy
$(\mathsf{R}_{1}^{1\rightarrow 2},\mathsf{R}_{2}^{1\rightarrow 2})$ $\simeq\!(\log_{2}(1\!+\!{\gamma_{1}r_{2}^{2}}/{(\gamma_{2}r_{1}^{2})}),\mathcal{O}(\log M)),$ (30a)
$(\mathsf{R}_{1}^{2\rightarrow 1},\mathsf{R}_{2}^{2\rightarrow 1})$ $\simeq\!(\mathcal{O}(\log M),\log_{2}(1\!+\!{\gamma_{2}r_{1}^{2}}/{(\gamma_{1}r_{2}^{2})})).$ (30b)
If UTs are positioned in different directions, we have
$(\mathsf{R}_{1}^{1\rightarrow 2},\mathsf{R}_{2}^{1\rightarrow 2})$ $\approx(\mathsf{R}_{1}^{2\rightarrow 1},\mathsf{R}_{2}^{2\rightarrow 1})$ (31a)
$\simeq(\mathcal{O}(\log M),\mathcal{O}(\log M)).$ (31b)
The above results indicate that the MAC capacity region under the FF model can expand unboundedly with the number of array elements, rendering it impractical. Additionally, when UTs face different directions, the FF MAC capacity region roughly forms a rectangle and is more extensive than that when UTs are in the same direction.
By comparing the NF capacity with its FF counterpart, we have the following observations:
  • Different from the FF MAC capacity that degenerates for co-directional UTs due to the significant IUI, the NF capacity can be preserved when UTs are located at different locations (different directions or distances), which underscores superior flexibility and robust interference management capabilities inherent to NFC. This additional resolution in the range domain enables space-division multiple access [28] from FF angle-division multiple access to NF range-division multiple access, demonstrating the NFC’s potential as a promising approach for efficient and interference-free communication systems.
III-C Comparison with Linear Combiners
[…]
${\mathsf{R}}_{k}=\log_{2}\bigg{(}1+\frac{\lvert{\mathbf{v}}_{k}^{\mathsf{H}}{\mathbf{h}}_{k}\rvert^{2}P_{k}}{\sigma^{2}\lVert{\mathbf{v}}_{k}\rVert^{2}+\sum_{k^{\prime}\neq k}\lvert{\mathbf{v}}_{k}^{\mathsf{H}}{\mathbf{h}}_{k^{\prime}}\rvert^{2}P_{k^{\prime}}}\bigg{)}.$ (32)
For the sake of brevity, we continue to consider the two-user case with $K=2$.
III-C1 Optimal Linear Combining
We begin with the optimal linear combiner, which is designed by first whitening the interference-plus-noise term and then applying a maximal-ratio combiner to the resulting effective channel. According to [27], the optimal linear combiner for UT $k$ is given by:
${\mathbf{v}}_{k}=(\sigma^{2}{\mathbf{I}}_{M}+P_{k^{\prime}}{\mathbf{h}}_{k^{\prime}}{\mathbf{h}}_{k^{\prime}}^{\mathsf{H}})^{-1}{\mathbf{h}}_{k},$ (33)
where $k^{\prime}\neq k$. Substituting (33) into (32) gives
${\mathsf{R}}_{k}=\log_{2}(1+P_{k}{\mathbf{h}}_{k}^{\mathsf{H}}(\sigma^{2}{\mathbf{I}}_{M}+P_{k^{\prime}}{\mathbf{h}}_{k^{\prime}}{\mathbf{h}}_{k^{\prime}}^{\mathsf{H}})^{-1}{\mathbf{h}}_{k}),$ (34)
which, together with the Woodbury matrix identity, yields
${\mathsf{R}}_{1}+{\mathsf{R}}_{2}=\sum_{k=1}^{2}\log_{2}\Big{(}1+\gamma_{k}{\mathsf{g}}_{k}\Big{(}1-\frac{\gamma_{k^{\prime}}{\mathsf{g}}_{k^{\prime}}}{1+\gamma_{k^{\prime}}{\mathsf{g}}_{k^{\prime}}}\rho\Big{)}\Big{)}\triangleq{\mathsf{R}}_{{\mathsf{opt}}}.$ (35)
III-C2 Maximal-Ratio Combining
Next, we consider maximal-ratio combining (MRC), which aims to maximize the channel gain for each UT. This yields $\mathbf{v}_{k}=\mathbf{h}_{k}$. Substituting $\mathbf{v}_{k}=\mathbf{h}_{k}$ into (32) and performing some basic mathematical manipulations, we express the sum-rate as follows:
${\mathsf{R}}_{1}+{\mathsf{R}}_{2}=\sum_{k=1}^{2}\log_{2}\Big{(}1+\gamma_{k}{\mathsf{g}}_{k}\Big{(}1-\frac{\gamma_{k^{\prime}}{\mathsf{g}}_{k^{\prime}}\rho}{1+\gamma_{k^{\prime}}{\mathsf{g}}_{k^{\prime}}\rho}\Big{)}\Big{)}\triangleq{\mathsf{R}}_{{\mathsf{mrc}}}.$ (36)
III-C3 Zero-Forcing Combining
Finally, we consider zero-forcing (ZF) combining, which aims to nullify the IUI. In this case, ${\mathbf{v}}_{k}^{\mathsf{H}}{\mathbf{h}}_{k^{\prime}}=0$ for $k^{\prime}\neq k$. To further maximize the channel gain of UT $k$, we set ${\mathbf{v}}_{k}$ as the projection of $\mathbf{h}_{k}$ onto the space orthogonal to $\mathbf{h}_{k^{\prime}}$, yielding [27]
${\mathbf{v}}_{k}=(\mathbf{I}_{M}-\mathbf{h}_{k^{\prime}}\mathbf{h}_{k^{\prime}}^{\mathsf{H}}/\|\mathbf{h}_{k^{\prime}}\|^{2})\mathbf{h}_{k}.$ (37)
Substituting (37) into (32) and performing some simple manipulations, we obtain
${\mathsf{R}}_{1}+{\mathsf{R}}_{2}=\sum_{k=1}^{2}\log_{2}(1+\gamma_{k}\mathsf{g}_{k}(1-\rho))\triangleq{\mathsf{R}}_{{\mathsf{zf}}}.$ (38)
III-C4 Summary and Discussion
Through the derivations above, we observe that the sum-rate achieved by the considered linear combining schemes can be expressed as a function of the channel gains and the CCF, i.e.,
${\mathsf{R}}_{\ell}=\sum_{k=1}^{2}\log_{2}(1+\gamma_{k}\mathsf{g}_{k}(1-f_{\ell}(\gamma_{k^{\prime}}\mathsf{g}_{k^{\prime}},\rho))),$ (39)
where $\ell\in\{{\mathsf{opt}},{\mathsf{mrc}},{\mathsf{zf}}\}$, $f_{{\mathsf{opt}}}(x,z)\triangleq\frac{xz}{1+x}$, $f_{{\mathsf{mrc}}}(x,z)\triangleq\frac{xz}{1+xz}$, and $f_{{\mathsf{zf}}}(x,z)\triangleq z$. By substituting (15) and (16) into (39), we obtain the NF sum-rate expressions under the linear combiners, which can be written as follows:
${\mathsf{R}}_{\ell}=\sum_{k=1}^{2}\log_{2}(1+\gamma_{k}\mathsf{g}_{k}(1-f_{\ell}(\gamma_{k^{\prime}}\mathsf{g}_{k^{\prime}}^{\mathsf{n}},\rho_{\mathsf{n}}))).$ (40)
Moreover, it can be observed that $\mathsf{C}_{\mathsf{MAC}}\geq{\mathsf{R}}_{{\mathsf{opt}}}\geq{\mathsf{R}}_{{\mathsf{zf}}}$ and $\mathsf{C}_{\mathsf{MAC}}\geq{\mathsf{R}}_{{\mathsf{opt}}}\geq{\mathsf{R}}_{{\mathsf{mrc}}}$. However, the relationship between ${\mathsf{R}}_{{\mathsf{mrc}}}$ and ${\mathsf{R}}_{{\mathsf{zf}}}$ depends on the values of $\rho$ and ${\mathbf{g}}_{k}$.
As discussed in Section III-A1, we have $\lim_{M_{x},M_{z}\rightarrow\infty}\rho_{\mathsf{n}}\approx 0$ under the NF channel model. This suggests that
$\lim\nolimits_{M_{x},M_{z}\rightarrow\infty}{\mathsf{R}}_{\ell}\approx\sum_{k=1}^{2}{\log_{2}\left(1+\gamma_{k}\mathsf{g}_{k}^{\mathsf{n}}\right)}$ (41)
holds for $\ell\in\{{\mathsf{opt}},{\mathsf{mrc}},{\mathsf{zf}}\}$.
@@ III-C4 Summary and Discussion
The above discussion implies that in NF MAC, when the number of array elements is sufficiently large, the considered linear combining schemes can approach the upper bound of the sum-rate capacity, similar to the capacity-achieving schemes. This occurs because the NF effect can make the UTs’ channels orthogonal to each other, resulting in zero IUI.
(a) $(\theta_{2},\phi_{2})=(\frac{\pi}{3},\frac{2\pi}{3})$.
(b) $(\theta_{2},\phi_{2})=(\frac{\pi}{3}\!+\!\frac{\pi}{60},\frac{2\pi}{3}\!+\!\frac{\pi}{60})$.
Figure 3: $\Gamma_{{\mathsf{ul}},\ell}$ versus $M$ for different combiners with $(\theta_{1},\phi_{1})=(\frac{\pi}{3},\frac{2\pi}{3})$ and the other parameter settings given in Section VI.
However, it is important to note that the conclusion in Remark 5 only holds when $M_{x},M_{z}\rightarrow\infty$. In practical NFC systems with large yet finite values of $M_{x},M_{z}$, there will be a gap between the achievable sum-rate and its upper bound $\sum_{k=1}^{2}\log_{2}(1+\gamma_{k}\mathsf{g}_{k}^{\mathsf{n}})$, which is evaluated as follows:
$\Gamma_{{\mathsf{ul}},\ell}\triangleq\frac{\sum\nolimits_{k=1}^{2}\log_{2}(1+\gamma_{k}\mathsf{g}_{k}^{\mathsf{n}}(1-f_{\ell}(\gamma_{k^{\prime}}\mathsf{g}_{k^{\prime}}^{\mathsf{n}},\rho_{\mathsf{n}})))}{\sum\nolimits_{k=1}^{2}{\log_{2}\left(1+\gamma_{k}\mathsf{g}_{k}^{\mathsf{n}}\right)}}.$ (42)
Due to the complexity involved in $\rho_{\mathsf{n}}$ (Eq. (16)), a quantitative comparison of $\Gamma_{{\mathsf{ul}},\ell}$ for different $\ell$ is challenging. Therefore, we use numerical results to explore the relationship between $\Gamma_{{\mathsf{ul}},\ell}$ and the antenna number $M$, as shown in Figure 3. For comparison, the ratio between the capacity achieved by SIC and the upper bound is also presented. As demonstrated, with the increasing number of antennas, all of the discussed combining schemes achieve a sum-rate close to their upper bound. Among the discussed linear combining schemes, the optimal linear combining scheme yields the best sum-rate performance. This indicates that the optimal combining scheme can leverage the weak IUI arising from the NF effect to the greatest extent among the three linear combining schemes. For the other two schemes, we observe that the MRC scheme is preferred over ZF combining with fewer antennas, while ZF combiner outperforms MRC combiner when the antenna number exceeds a certain threshold, demonstrating that eliminating interference is more beneficial than maximizing channel gain in larger antenna arrays. Furthermore, the sum-rates achieved by linear combiners are lower than the capacity when $M$ is small, but nearly the same as the capacity after $M$ reaching about $10^{5}$, which suggests that in NFC, linear combiners are able to provide adequate performance compared to optimal SIC for large-scale arrays.